OpenStudy (anonymous):

Write the equation of the line in Slope-Intercept form, Then graph the equation by using either the slope-intercept equation or by using x- and y-intercepts. The line contains points (3, 5) and (3, 8).

OpenStudy (anonymous):

I have already did this problem, just need to make sure I have done it right. 0 is the slope obviously because the line is always on the x-axis.

OpenStudy (anonymous):

Slope is undefined, which means its a vertical line. So the equation is gonna be x=whatever the x coordinate always is. In this case x=3

OpenStudy (anonymous):

Ahh, okay. Thanks, that's all I needed.

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

I have one more problem like this. Can you help me through it?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Alright, thanks. Here it is. The line contains points (2, -1) and (0, 4).

OpenStudy (anonymous):

And you want slope intercept form?

OpenStudy (anonymous):

Yes, same as the others.

OpenStudy (anonymous):

Alright, well if you notice, one of your points is the y intercept. So if we find the slope we can just plug those in and were done.

OpenStudy (anonymous):

Slope is -2

OpenStudy (anonymous):

You should have got -5/2

OpenStudy (anonymous):

Hmmm, trying again one min.

OpenStudy (anonymous):

Ahh, okay. Gothcha.

OpenStudy (anonymous):

Alright, slope intercept form is y=mx+b where m=slope and b=y coordinate of the y intercept. So plug those in and what do you get?

OpenStudy (anonymous):

2 = -5/2-1 + 2 or something like that....

OpenStudy (anonymous):

No, all you have to plug in is m and b. y=-5/2x+4

OpenStudy (anonymous):

Oh...yeah. heh I'm always gettin the steps mixed up.

OpenStudy (anonymous):

Yep. So now you're done and thats your answer. Understand how to do it now?

OpenStudy (anonymous):

Well yes and no. I'll most likely forget anyway.